Lee Holmes deposited $17,000 in a new savings account at 9% interest compounded semiannually. At the beginning of year 4, Lee deposits an additional $42,000 at 9% interest compounded semiannually. At the end of 6 years, what is the balance in Lee’s account? (Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1

Period of deposit of $17,000 is 6 years and that of $42,000 is 3 years (beginning of year 4 to end of year 6). Interest rate is 9% compounded semi annually.

Balance at the end 6 years is the sum of future values of both the deposits.

FV of $17,000 = $28,829.98 calculated as follows:

B C D E 1 Future Value 2 7 3 Future value of an amount is calculated using the formula FV=P(1+r)^n 4 Where P= Principal (Pres

FV of $42,000 = $54,694.93 calculated as follows:

А B C D 1 Future Value 2 3 Future value of an amount is calculated using the formula FV=P(1+r)^n 4 Where P= Principal (Presen

Balance at the end 6 years (sum of future values of both the deposits)= $28,829.98 + $54,694.93

= $ 83,524.91